OpenStudy (brittany):

solve using quadratic formula 3x^2+12x+8=0

OpenStudy (brittany):

a=3 b=12 c=8 -12±{12^2-4(3)(8)}/2(3) -12±{144-96}/2(3) -12±{48}/6

OpenStudy (brittany):

i used{} to represent square root symbol. im not sure if i did it right cos i cannot finihs it

OpenStudy (anonymous):

whats the final answer?

OpenStudy (brittany):

im not sure because their is no square root of 48?

OpenStudy (anonymous):

\[(-12+4\sqrt{3})/6\] and \[(-12-4\sqrt{3})/6\]

OpenStudy (brittany):

4 square roots of thre is equal to 48?

OpenStudy (brittany):

how do i solver for those next two equations then?

OpenStudy (anonymous):

finally you will get \[(-6+2\sqrt{3})/3\] and \[(-6-2\sqrt{3})/3\]

OpenStudy (anonymous):

yup\[4\sqrt{3}=\sqrt{48}\]

OpenStudy (brittany):

(−6+2√3)/3=(−2+2√3)/1 (−6−2√3)/3=(−2−2√3)/1

OpenStudy (anonymous):

nope

OpenStudy (brittany):

well hwat should i have done

OpenStudy (anonymous):

Sorry I was at lunch. Let me take a look

OpenStudy (anonymous):

Ok starting from what you had here: \[\frac{-12 \pm \sqrt{48}}{6}\]\[=-\frac{12}{6} \pm \frac{4\sqrt{3}}{6}\]\[=-2 \pm \frac{2\sqrt{3}}{3}\]

OpenStudy (brittany):

ok

OpenStudy (anonymous):

So I find it is often easier to break the fraction up into two then simplify each part

OpenStudy (brittany):

ok

OpenStudy (brittany):

hello?:(

OpenStudy (anonymous):

hello what? Did you have another question?

OpenStudy (brittany):

how do i finish the problem?

OpenStudy (anonymous):

polpak's answer is better than mine

OpenStudy (brittany):

whats the answer? im so confused

OpenStudy (anonymous):

both are correct

OpenStudy (brittany):

those arent the answers thoguh usually its x= # or #

OpenStudy (anonymous):

One is the + version. One is the - version.

OpenStudy (anonymous):

\[-2+(2\sqrt{3}/3)\] and \[-2-(2\sqrt{3}/3)\]

OpenStudy (anonymous):

\[x = -2 + \frac{2\sqrt{3}}{3}\]or\[x = -2 - \frac{2\sqrt{3}}{3}\] That's what\[x=-2\pm \frac{2\sqrt{3}}{3}\] means

OpenStudy (brittany):

i entered x= -2+2√3/3 or -2-2√3/3 and it was incorrect

OpenStudy (anonymous):

Dunno what to tell you, those are the right answers. Could also look like: \[\frac{2}{3}(-3 +\sqrt{3})\] and \[\frac{2}{3}(-3 -\sqrt{3})\]

OpenStudy (brittany):

hmm ok well thanks anyway i guess